the board of directors of a coporation grant a bonus of $5,000 to divided between two employees. the senior employee receives $1200 more then the junior employee. how much does each employee receive?
Answer by WannabeCAgirl83(35)


junior employee = x
senior employee = x+1200

Junior employee and senior employee together receive 5000$:
x + x+1200 = 5000

x + x+1200 = 5000 | -1200 on both sides
x + x = 3800 | re-arranging
2x = 3800 | :2 on both sides
x = 1900

junior employee = 1900
senior employee = 3100 (1900+1200)